OpenStudy (anonymous):

Can someone explain how I could simplify this, STEP by STEP?: 2sqrt3 / sqrt5 + sqrt2

OpenStudy (anonymous):

\[2\sqrt{3} \div \sqrt{5} + \sqrt{2}\]

OpenStudy (callisto):

is the question like this?